OpenStudy (aaronandyson):

The ages of two sisters are 11 years and 14 years.In how many years time will the product of their age be 304?

OpenStudy (aaronandyson):

@mathrulezz

OpenStudy (welshfella):

(11 + x)(14 + x) = 304

OpenStudy (welshfella):

154 + 25x + x^2 = 304 x^2 + 25x - 150 = 0 solve for x

OpenStudy (aaronandyson):

10 & 15?

OpenStudy (aaronandyson):

I got 5